OpenStudy (anonymous):

Which is 625 written as a power of a number? A. 3^4 B. 3^5 C. 5^3 D. 5^4

OpenStudy (anonymous):

625 ends in 5 so it can't be a or b. see what the others give. @hbaker1 please read the code of conduct. don't just give out answers, let people work and discover.

OpenStudy (anonymous):

sorry, you kept complaining about my posts when i was trying to help someone learn and you didn't want to receive the notices. you can remove your posts in that situation and you'll not receive any notices.

OpenStudy (anonymous):

A. 3x3x3x3=81

OpenStudy (math&ing001):

625 = 25 * 25 = 5² * 5² = 5^4